# PEER: Expertizing Domain-Specific Tasks with a Multi-Agent Framework and Tuning Methods

### v1

@@ 1 Introduction
The main contributions of this study include: 1.Providing and open-sourcing the PEER [^1] https://github.com/alipay/agentUniverse framework, characterized by its conciseness, effectiveness, and cost-efficiency, for effectively tackling domain-specific tasks. In experiments, it demonstrates superior performance compared to BabyAGI.
2.Proposing a customized agent tuning strategy for 10-billion-parameter models, achieving performance comparable to GPT-4.
3.Constructing and open-sourcing a dataset for use within the PEER framework, applicable to agent training, pre-training, and supervised fine-tuning in various financial analysis scenarios.
@@ 4.1 PEER Framework experiment
Dataset Since the main usage scenario of PEER framework is the interpretation and analysis of domain events and problems, we mainly tested and compared the performance of PEER on the dataset of financial QA. We sampled hundreds of professional questions [^3] https://github.com/alipay/agentUniverse/tree/master/datasetfrom our business scenarios and divided them into nine categories. Details of the dataset distribution are shown in Table 4.

### v2

@@ Abstract
footnotetext: 1These authors contributed equally to this work.footnotetext: Corresponding author: jihan.hanji@antgroup.com
@@ 1 Introduction
The main contributions of this study include: 1. Providing and open-sourcing the PEER***https://github.com/alipay/agentUniverse framework, characterized by its conciseness, effectiveness, and cost-efficiency, for effectively tackling domain-specific tasks. In experiments, it demonstrates superior performance compared to BabyAGI.
2. Proposing a customized agent tuning strategy for 10-billion-parameter models, achieving performance comparable to GPT-4.
3. Constructing and open-sourcing a dataset for use within the PEER framework, applicable to agent training, pre-training, and supervised fine-tuning in various financial analysis scenarios.
@@ 4.1 PEER Framework experiment
Dataset Since the main usage scenario of PEER framework is the interpretation and analysis of domain events and problems, we mainly tested and compared the performance of PEER on the dataset of financial QA. We sampled hundreds of professional questionshttps://github.com/alipay/agentUniverse/tree/master/datasetfrom our business scenarios and divided them into nine categories. Details of the dataset distribution are shown in Table 4.